the sum of eight and twice a number is equal to 2 less than the number. whats the number.
Found 2 solutions by richwmiller, checkley77:
Answer by richwmiller(17219)

sum means add
eight means 8
twice a number means 2n
is equal means =
2less means -2
the number means n
8+2n=-2+n
8+2n=n-2
add 2 to both sides (bs)
10+2n=n
subtract 2n from bs
10=-n
n=-10
8+*2(-10)=-10-2
8-20=-12
-12=-12

Answer by checkley77(12844)

8+2x=x-2
2x-x=-2-8
x=-10 ans.
Proof:
8+2*-10=-10-2
8-20=-12
-12=-12